OpenStudy (anonymous):

Find the sum and product of the roots. 3x 2 - 4x - 7 = 0

OpenStudy (anonymous):

a. sum = -4/7, product = -7/3 b. sum = 4/3 product = -7/3 c. sum = -4, product = -7 d. sum = 3/5 , product =7/3

OpenStudy (tkhunny):

Divide by 3 \(x^{2} - (4/3)x - 7/3 = 0\) Ponder this: \((x-a)(x-b) = 0\) \(x^{2} - (a+b)x + ab = 0\)